At what quoted annual interest rate must $159,000 be invested so that it will grow to be $467,000 in 10 years if interest is compounded weekly?

11.38%
10.79%
0.21%
8.77%
15.47%
Answer #1

We use the formula:
A=P(1+r/52)^52n
where
A=future value
P=present value
r=rate of interest
n=time period.

467000=159000*(1+r/52)^(52*10)

(467000/159000)^(1/520)=(1+r/52)

(1+r/52)=1.002074119

r=(1.002074119-1)*52

=10.79%(Approx).